OpenStudy (anonymous):

The number of people in a town of 10,000 who have heard a rumor started by a small group of people is given by the following function: a. How many people were in the group that started the rumor? b. How many people have heard the rumor after 1 day? After 5 days? c. How long will it be until 1,000 of the people in the town have heard the rumor? WILL GIVE MEDAL!!! Need help asap though!

OpenStudy (shamil98):

Where is the "following function"?..

OpenStudy (anonymous):

N(t) = 10,000/ 5 + 1245e ^ -0.97t

OpenStudy (anonymous):

that is the function :)

OpenStudy (shamil98):

Okay. Let's start with A. when t = 0 .

OpenStudy (shamil98):

evaluate it for when t = 0 you get. n(0) = 10000/(5+1245(e^-0.97(0) = 10000/1250 = 8 so 8 people.

OpenStudy (shamil98):

All you have to do is substitute for the value of t , and evaluate the functions. For b.

OpenStudy (anonymous):

ok

OpenStudy (shamil98):

Now, for C. You want to know when 1000 people have heard the rumor..so substitute N with 1000 and solve for t.

OpenStudy (shamil98):

1000 = 10000/(5+1245e^-.97t) 1/10 = 1/(5+1245e^-.97t) 5 + 1245e^-.97t = 10 1245e^.97t = 5 divide both sides.. e^-.97t =.0040 take the natural log of both sides. -.97t = ln .004 simplify. t = 5.7

OpenStudy (anonymous):

about 5 days ok so 5.7 weeks?

OpenStudy (anonymous):

and for b I get 2471 for 1 day and 4359.79 for b

OpenStudy (shamil98):

t = 5.7 so about 6 days.

OpenStudy (shamil98):

after 1 day 21 people will hear the rumor, after 5 days bout 678 people will hear the rumor...

OpenStudy (anonymous):

for 5 days day I get 678 and and 21 for 1 day

OpenStudy (anonymous):

right sorry I made an error in my calculator

OpenStudy (shamil98):

yeah. good job :)

OpenStudy (anonymous):

thanks can you help me on one more problem?

OpenStudy (anonymous):

Solve each equation for x: 5 = log3(x^2 + 18) and -2 = log(3x+5)

OpenStudy (shamil98):

5 = log 3(x^2+18) 10^-5 = 3(x^2+18) does that help?

OpenStudy (shamil98):

-2 = log (3x+5) 10^-2 = 3x + 5 just solve for x now :)

OpenStudy (anonymous):

ok I think I have the answers could you post both for me to double check?

OpenStudy (shamil98):

I don't have the first one done, too lazy to do a quadratic.. but for the second one x = -1.663

OpenStudy (anonymous):

ok I got that one too :)

OpenStudy (anonymous):

can u answer this last one?

OpenStudy (anonymous):

e^x e^(x+1) = 1

OpenStudy (shamil98):

\[e^x e^{x+1}=1\] not sure how to do this one sorry.

OpenStudy (shamil98):

but im pretty sure it involves using the natural log >.<

OpenStudy (anonymous):

ok thanks